OpenStudy (jacksparr0w):

plz help. willl fan u and metal u 9 x 8 = a

OpenStudy (turtwig):

do you know what 9x8 is?

OpenStudy (smart.420.smoker):

72

OpenStudy (jacksparr0w):

I think its times

OpenStudy (cas_fangirl_14):

a=72

OpenStudy (turtwig):

right, so 9 times 8 equals 72. a=72

OpenStudy (jacksparr0w):

thank u both! but i was only abel to give 1 medal

OpenStudy (jacksparr0w):

thank u turtwig

OpenStudy (cas_fangirl_14):

nah its cool

OpenStudy (ericamarie7):

oh wow lmao

OpenStudy (jacksparr0w):

wat is it erica?

OpenStudy (turtwig):

i just want you to learn this stuff, it's pretty easy if you know what you are doing :)

OpenStudy (jacksparr0w):

ok thank u @turtwig i still have more

OpenStudy (frozenghost127):

suck 1

OpenStudy (shadowlegendx):

Here's a good trick to solving multiplication problems having to do with 9. 9 times 8. Go 1 value down on the number 9 is being multiplied by, which is 7. The number that adds up with that, to get 9, is your second number. Examples: 9 times 8 8-1 = 7 7 + 2 = 9 9 times 8 = 72

OpenStudy (shadowlegendx):

9 times 5 5-1 = 4 4 + 5 = 9 9 times 5 = 45

OpenStudy (cas_fangirl_14):

Please tell me you are in 3rd grade or something @JackSparr0w

OpenStudy (shadowlegendx):

9 times 7 7-1 = 6 6 +3 = 9 9 times 7 = 63

OpenStudy (jacksparr0w):

oh wow thanks shadowlegend i will right that down bc its really helpful

OpenStudy (shadowlegendx):

Only works up to 9 times 9

OpenStudy (shadowlegendx):

No problem

OpenStudy (cas_fangirl_14):

9 18 27 36 45 54 63 72 81 See the pattern?